OpenStudy (anonymous):

Two gears are connected and are rotating simultaneously. The smaller gear has a radius of 4 inches, and the larger gear has a radius of 7 inches. two circles touching at one point. Larger circle has radius of 7 inches. Smaller circle has radius of 4 inches. Part 1: What is the angle measure, in degrees and rounded to the nearest tenth, through which the larger gear has rotated when the smaller gear has made one complete rotation? Part 2: How many rotations will the smaller gear make during one complete rotation of the larger gear? Show all work.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me think

jimthompson5910 (jim_thompson5910):

we basically have this picture going on 

jimthompson5910 (jim_thompson5910):

what is the circumference of circle A?

OpenStudy (anonymous):

25.12

jimthompson5910 (jim_thompson5910):

let's leave it in terms of pi

jimthompson5910 (jim_thompson5910):

you should get 8pi

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

how about the larger circle?

OpenStudy (anonymous):

14pi

jimthompson5910 (jim_thompson5910):

if we divide 14pi by 8pi, we get (14pi)/(8pi) = 14/8 = 7/4

jimthompson5910 (jim_thompson5910):

convert 7/4 to a decimal to get 7/4 = 1.75 basically this number tells us that the smaller circle will do 1.75 full rotations as it rolls around the larger circle. Basically the smaller circle does 1.75 rotations per 1 rotation of the larger circle

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

take this ratio 1.75 rotations of A: 1 rotation of B and divide both parts by 1.75 to get (1.75/1.75) rotation of A: (1/1.75) rotation of B 1 rotation of A: 0.57142857 rotation of B

jimthompson5910 (jim_thompson5910):

so 1 full revolution of circle A means circle B only does approx 0.57142857 of a full revolution

OpenStudy (anonymous):

so how to answer part 1

jimthompson5910 (jim_thompson5910):

well you simply multiply 0.57142857 by the number of degrees in a full rotation

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

part 2 is pretty much answered when I said how many times A revolves as it rolls around B